. Imagine you want to determine the number of genes that influence the color pattern in snakes. You cross two inbred (i.e. homozygous for all genes) snakes with different color patterns, then intercross the F1s. At the point where you have produced 972 F2 offspring, you have one offspring that has a color pattern that is the same as each of the parents (offspring number 478 resembles the father, and offspring number 684 resembles the mother). How many genes influence skin color in the snake?
Answer #1

This is the case of polygenuc inheritance. In this case, there is found involvement of more than two genes.

There is found three types of offspring in F2:

* One offspring have both parents color

* 478 have father type color.

* 684 have mother's type color.

That's why there is found 3 phenotypes.

Number of genes = n

We know,

Number of phenotype = (2× number of genes)+1

3 = 2×n +1

3-1= 2×n

2/2 = n

n= 1

That's why in this case, one gene encodes the snake color pattern.